Use the Taylor series expansion to find the general solution of each of the following

2x2y''-3xy'+(x+1)y=0

y(1)=4 y'(1)=1


1
Expert's answer
2021-08-12T17:29:05-0400

The Taylor series expansion of the function "f(x)" has the form: "f(x)=f(a)+f'(a)(x-a)+\frac{f''(a)}{2!}(x-a)^2+\frac{f'''(a)}{3!}(x-a)^3+..."

In our case the point is "a=1". We receive: "f(x)=f(1)+f'(1)(x-1)+\frac{f''(1)}{2!}(x-1)^2+\frac{f'''(1)}{3!}(x-1)^3+..."

We take derivatives and get:

"f'(x)=f'(1)+f''(1)(x-1)+\frac{f'''(1)}{2!}(x-1)^2+..."

"f''(x)=f''(1)+f'''(1)(x-1)+..."

The initial conditions take the form: "f(1)=4", "f'(1)=1". We substitute the expansion and receive:

"2x^2(f''(1)+f'''(1)(x-1)+...)-3x(f'(1)+f''(1)(x-1)+\frac{f'''(1)}{2!}(x-1)^2+...)+(x+1)(f(1)+f'(1)(x-1)+\frac{f''(1)}{2!}(x-1)^2+\frac{f'''(1)}{3!}(x-1)^3+...)=0"

We denote: "z=x-1" and receive: "2(z+1)^2(f''(1)+f'''(1)z+...)-3(z+1)(f'(1)+f''(1)z+\frac{f'''(1)}{2!}z^2+...)+(z+2)(f(1)+f'(1)z+\frac{f''(1)}{2!}z^2+\frac{f'''(1)}{3!}z^3+...)=0"

We rewrite and receive:

"2(z^2+2z+1)(f''(1)+f'''(1)z+...)-3(z+1)(f'(1)+f''(1)z+\frac{f'''(1)}{2!}z^2+...)+(z+2)(f(1)+f'(1)z+\frac{f''(1)}{2!}z^2+\frac{f'''(1)}{3!}z^3+...)=0"

We collect the terms near "1,z,z^2,z^3,..." and get: "2f''(1)-3f'(1)+2f(1)+z(4f''(1)+2f'''(1)-3f''(1)-3f'(1)+f(1)+2f'(1))+..."

We set: "f(1)=4" and "f'(1)=1" and get: "2f''(1)-3+8=0". It implies "f''(1)=-2.5". The term near "z" yields:"-10+2f'''(1)-7.5-3+4+2=2f'''(1)-20.5+6=-14.5". Thus, "f'''(1)=-7.25". We obtained first terms of the expansion: "f(1)=4", "f'(1)=1", "f''(1)=-2.5", "f'''(1)=-14.5". In case we calculate the coefficients near "z^4,z^5" and other terms, we will receive all the other derivatives ("f^{(4)}(1),f^{(5)}(1),..." ) and the explicit form of "f(x)".